The Day garden has 24 students in its 3-and 4-year old program. There are twice as many 4-yr olds as there are 3-year olds. How many of each age are there?
Answer by TimothyLamb(4379)

x = number of 3 year olds
y = number of 4 year olds
---
x + y = 24
y = 2x
---
put the system of linear equations into standard form
---
x + y = 24
2x - y = 0
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = number of 3 year olds = 8
y = number of 4 year olds = 16
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php